OpenStudy (jazzyluvsu123):

change 2&3/100 to percent HELP PLEASE show work FREE MEDAL !!!

OpenStudy (anonymous):

you see, 2$3 isn't a number :DDD

OpenStudy (jazzyluvsu123):

I meant 2&3/100 @sourwing sorry

OpenStudy (anonymous):

what's 2&3? lol

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\%\ =\cfrac{{\color{red}{ a}}}{100}\\ \quad \\ \quad \\ {\color{red}{ 2\frac{3}{100}}}\%=\cfrac{{\color{red}{ 2\frac{3}{100}}}}{100}\)

OpenStudy (jdoe0001):

keep in mind that \(\bf 2\cfrac{3}{100}\implies \cfrac{2\cdot 100+3}{100}\)

OpenStudy (jazzyluvsu123):

omg @sourwing \[2 \frac{ 3 }{ 100 }\]

OpenStudy (jazzyluvsu123):

would the answer be 203% @jdoe0001

OpenStudy (jdoe0001):

so....what would get for the fraction for \(\bf \bf 2\cfrac{3}{100}\implies \cfrac{2\cdot 100+3}{100}\quad ?\)

OpenStudy (jazzyluvsu123):

huh? @jdoe0001

OpenStudy (jdoe0001):

you need to change the fraction, first, from Mixed to just having 1 numerator and denominator.... thus

OpenStudy (jdoe0001):

hmmm oh.... wait.... I got it backwards you have the decimal and want the %... ok

OpenStudy (jazzyluvsu123):

is 203% right ?

OpenStudy (jdoe0001):

hmmm how did you get that?

OpenStudy (jazzyluvsu123):

\[divide 3\div100 put whole number then decimal and 203%\]

OpenStudy (jazzyluvsu123):

I divided 3/100 out 2 down then put . and divided from there @jdoe0001

OpenStudy (jdoe0001):

\(\bf 2\frac{3}{100}\cdot 100\implies \cfrac{2\cdot 100+3}{100}\cdot 100\implies \cfrac{203}{\cancel{100}}\cdot \cancel{100}\)

OpenStudy (jazzyluvsu123):

soo.. was my answer right I really need to move on

OpenStudy (jdoe0001):

yes

OpenStudy (jazzyluvsu123):

thanks @jdoe0001 Free medal for u

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

ah i see XDD